# Self-organisation - the underlying principle and a general model

### v2

@@ V. Extension of the modelling framework
As the DOFs are usually continuous, the general canonical partition function is then
@@
VI. Example application of the formalism
To illustrate the application of this formalism, consider the phenomenon of self-organised laning He71 ; HeMo95 ; Chetal00 . Laning is a spontaneous formation of lanes as agents (pedestrians, cars, etc.) move in crowded spaces. It is the result of agents aiming to move in a particular direction or reach a particular destination without bumping too often and too violently into others. While collisions, and their avoidance, are local, ordered system-wide laning often emerges. This SO makes the flow of agents more efficient, e.g., by minimising the time spent in the crowd before reaching the destination. The dynamic lanes pattern is but a minute fraction of all the possible ways that the agents can move. In a simple version of this problem, $N$ agents move in one direction along a long strip of width that allows only few agents to move side by side. They can move at one of two speeds $v_{1}$ and $v_{2}>v_{1}$ and at any angle relative to the forward direction, $\pi/2<\theta<\pi/2$. Moving at $\theta=0$ is fastest, but can be hampered by collisions with more slowly-moving agents. Attempting to overtake those by moving sideways may incur collisions with other agents. Here, the agents are the quasi-particles and q=1,2,,Nq=1,2,...,N and each agent has two DOFs,the speed and direction of movement. For each agent, the most stable mode of movement is by avoiding collisions, achievable with the stability measure,
@@ VI. Example application of the formalism
where $q^{\prime}\Rightarrow q$ denotes the indices of the agents nearest to $q$ (identifiable, e.g., using Voronoi tessellation at each moment). The objective of reaching the destination efficiently is a constraint that also needs to be quantified in the survivability function and this is achievable by encouraging movement in the forward direction. Thus, a natural survivability function is
$F_{q}=J\sin^{2}{\theta_{q}}+\lambda\sum\limits_{q^{\prime}\Rightarrow q}\left(\vec{v}_{q}-\vec{v}_{q^{\prime}}\right)^{2}\ .$ (12)
@@
Further constraints and specifications can be readily introduced by adding appropriate terms to the survivability function. The corresponding partition function is then
$Z=$ $\int e^{-\sum\limits_{q=1}^{N}\left[J\sin^{2}{\theta_{q}}+\lambda\sum\limits_{q^{\prime}\Rightarrow q}\left(\vec{v}_{q}-\vec{v}_{q^{\prime}}\right)^{2}\right]}$
$\times\prod\limits_{q=1}^{N}\Big{\{}P\left(v_{q},\theta_{q}\right)dv_{q}d\theta_{q}\Big{\}}\ ,$ (13)
@@
with, say, $P\left(v_{q},\theta_{q}\right)=\left[p\delta(v-v_{1})+(1-p)\delta(v-v_{2})\right]/\pi$ the density of states. It is clear that, by increasing $J$ and $\lambda$, the stable system-wide configurations are in a small subset that approximate two lanes, one moving at $v_{1}$ and the other at $v_{2}$. This partition function can be readily evaluated, but, as the aim here is only to illustrate the application of the general formalism, this is downstream from the current work.
VII. Concluding discussion
The principle underlying emergence of SO in out-of-equilibrium driven disordered systems has been formulated. It states that SO is the result of survivability of a minute subset of system configurations, or microstates, that stand out in their exceptional stability against the noise introduced by the driving. This principle is clearly supported by observations of SO in 2D granular systems, which evolve toward stable configurations of coordinated cell structures and stresses. Based on this principle, a statistical mechanical formalism has been constructed for SO in those 2D systems and its use outlined. Similar stability measures have been identified in 3D granular systems, which makes possible a straightforward extension to a similar model and its outline has been discussed in detail.
The principle goes beyond granular systems - driven disordered systems always experience noisy dynamics and, when their steady states self-organise, they exclude the overwhelming majority of theoretically possible configurations which do not survive the noise. It also goes beyond the previously proposed principle of minimisation of dissipation energy Cyetal21 ; Keetal23 , which is only a special case that applies to systems where energy is the main driver of the SO. The generality of this principle makes possible an extension of the statistical mechanics approach to describe emergence of SO in wider range of systems. This extension has been formulated here explicitly and it involves identification of quasi-particles; their DOFs; stability measures, which depend on the DOFs; and a survivability function, whose value decreases with the stability measures. The latter is used in a Boltzmann-like exponential, which determines the occurrence probability of system configurations, and a partition function has been constructed. From the partition function, all expectation values of the self-organised state can be derived. This complements the idea in Keetal23 of the system exploring the configuration space in that it parks on configurations that minimise locally the survivability function. The application of the general formalism has been illustrated by constructing a statistical mechanics model for a very simple version of the laning problem.
@@ VII. Concluding discussion
There are parallels between this formulation and its analogue in thermal systems, in which the microstates are the system-wide configurations and the particles’ energies and potentials are used as stability measures. The thermal survivability function is the free energy, which depends on the quasi-particles’ DOFs and whose lowering increases a microstate’s occurrence probability and its survivability. Following this understanding, some equilibrium macrostates can also be viewed as self-organised. An example is solidification under temperature lowering, which is analogous to increasing $J_{i}$ in (4). The overwhelming number of microstates with very energetic particles are strongly unstable and only the minute crystalline microstates survive to be observed. There are, however, differences. The energy in thermal systems is a constant of the motion, from which dynamic equations for the DOFs can be derived. The survivability function only describes stability in out-of-equilibrium systems and, as far as this author knows, cannot play such a role. Additionally, thermal systems incur indigenous thermal fluctuations while in a-thermal systems it is the driving that gives rise to noise. Nevertheless, the similarities to traditional statistical mechanics are useful as guidelines for constructing general models of SO.
One can argue that the above principle also applies to the SO that is evident in the evolution of biological species. Those systems are driven by many constraints: resources, predators, environmental, etc., which they must adapt to and survive. The dynamics are then extremely noisy because the constraints’ magnitudes and even number can fluctuate wildly. The principle of the survival of the fittest is exactly the same - only elements that can survive these fluctuations, be it cells or larger organisms, remain. While this may mean that, in principle, it could be possible to formulate a statistical mechanical model for species evolution, it is a daunting task due to the enormous number of potential DOFs and stability measures.
@@ Acknowledgement:
Discussions with J.L. England and D.I. Goldman are gratefully acknowledged.

### v3

@@ V. A general modelling framework
As the DOFs are usually continuous, the general canonical partition function is
@@
VI. Example application - Laning SO
To illustrate the application of the formalism, consider the phenomenon of self-organised laning He71 ; HeMo95 ; Chetal00 . Laning is a spontaneous formation of lanes as agents (pedestrians, cars, etc.) move in crowded spaces. It is the result of agents aiming to move in a particular direction or reach a particular destination without bumping too often and too violently into others. While collisions, and their avoidance, are local, ordered system-wide laning often emerges. This SO makes the flow of agents more efficient, e.g., by minimising the time spent in the crowd before reaching the destination. The global lanes pattern consists a minute fraction of all the possible ways that the agents can move. As the aim is to illustrate the formalism, rather than improve existing models of this phenomenon, simplifications are made for calculating the partition function proposed below.
In a simple version of this problem, $N$ agents move in one direction, say $y$, along a long strip of width that allows only few agents to move side by side. They can move at one of two speeds $v_{1}$ and $v_{2}>v_{1}$ and at any angle relative to the forward direction, $\pi/2<\theta<\pi/2$. Moving at $\theta=0$ is fastest, but can be hampered by collisions with more slowly-moving agents. Attempting to overtake those by moving sideways may incur collisions with other agents. Here, the agents are the quasi-particles and q=1,2,,Nq=1,2,...,N and each agent has two DOFs - the speed and direction of motion. An agent’s most sustainable mode of progress is by avoiding collisions, achievable with the stability measure,
@@ VI. Example application - Laning SO
where $q^{\prime}\Rightarrow q$ denotes the indices of the agents nearest to $q$ (identifiable, e.g., using Voronoi tessellation at each configuration). The objective of reaching the destination efficiently is a constraint that can be quantified in the survivability function by encouraging motion in the forward direction. A natural survivability function is
$F_{q}=J\tan^{2}{\theta_{q}}+\lambda\sum\limits_{q^{\prime}\Rightarrow q}\left[\left(\vec{v}_{q}-\vec{v}_{q^{\prime}}\right)^{2}+\left(v_{qy}-u_{max}\right)^{2}\right]\ .$ (14)
@@
Here, $J$ biases motion forward and the last term encourages motion forward as quickly as possible. Further constraints and specifications can be introduced by adding appropriate terms in $F_{q}$. The partition function is
$Z=\int e^{-\sum\limits_{q=1}^{N}F_{q}}\times\prod\limits_{q=1}^{N}\Big\{P\left(v_{q},\theta_{q}\right)dv_{q}d\theta_{q}\Big\}\ ,$
@@
with, say, $P\left(v_{q},\theta_{q}\right)=\left[p\delta(v-v_{1})+(1-p)\delta(v-v_{2})\right]/\pi$ the density of states. It is clear that the interplay between $J$ and $\lambda$ determines that the stable system-wide configurations are the small subset of those that approximate two lanes, one moving at $v_{1}$ and the other at $v_{2}$. The partition function allows us to predict the characteristics of these configurations and it has been evaluated explicitly in the supplemental material SM in an effective medium approximation for a somewhat more general case (eq. (27) there). Using it, the mean lane speed has been calculated
$\langle v\rangle=\frac{\sum\limits_{m=1}^{2}\frac{S_{m}}{\sqrt{\xi_{m}}}v_{m}e^{-2\lambda Kv_{m}\left(v_{m}-u_{y}-u_{max}\right)+\xi_{m}C_{m}^{2}}}{\sum\limits_{m=1}^{2}\frac{S_{m}}{\sqrt{\xi_{m}}}e^{-2\lambda Kv_{m}\left(v_{m}-u_{y}-u_{max}\right)+\xi_{m}C_{m}^{2}}}\ ,$ (15)
with the variables defined in the supplemental material. A plot of this expectation value is shown in Fig. 1, in which $T=1/\lambda$ is used as a scaled measure of the noise, in analogy to temperature in thermal systems. In very noisy systems, i.e., with many collisions between agents, the mean progress speed converges to the average of the lanes’ speeds, $\langle v\rangle\to\left(v_{1}+v_{2}\right)/2$. At very low noise, $T\to 0$, the speed is dominated by the fast movers, $\langle v\rangle\to v_{2}>v_{1}$. Since collisions rate (=noise) increases with agents density, this behaviour is consistent with the results in Heetal06 ; Moetal12 . This model also reveals an intriguing phase-like transition: while at high values of the bias parameter, $J$, $\langle v\rangle$ decreases monotonically with $T$, as $J$ decreases, a minimum merges at some $T_{c}$. This demonstrates that the approach presented here can yield rich behaviour even for simple models. Analysing this transition in more detail is downstream from the main thrust of this paper.
Figure 1: The mean speed of the two lanes, derived in the mean field approximation, as a function of the noise parameter, $T=1/\lambda$, for several values of the bias parameter $J$. A transition from a monotonic decay to a minimum at some $T_{c}$ emerges as $J$ decreases below $J_{c}\approx 1.33$.
VII. Relevance to biological SO
Both the general principle and statistical mechanics method, proposed here, can apply to biological systems: proteins, cells, organisms, and entire species. In these contexts, the principle translates straightforwardly to ‘the fittest survive’. The formalism can also be extended by inspecting the equivalence between inanimate and biological systems. An organism’s survival is based on functionalities and skills that enhance its ’longevity’ and, consequently, that of the species. The equivalent of noise and fluctuations are factors that could shorten longevity and survivability, e.g., genetic variability, intra-species competition for reproduction, availability of resources, existence of predators, environmental cataclysms, etc.. An individual biological system is the equivalent of a quasi-particle, whose configuration is equivalent to the set of biologically inherited and adapted functionalities and skills. Better such skills improve ’fitness’ and survivability equivalently to better stability measures that increase the lifetime of a configuration to withstand the disruptive noise.
The main difference is that biological systems are active and can adapt, evolve, and develop feedback mechanisms to modify specific functionalities. This makes it possible for these systems to respond to noise in ways that passive systems cannot. In other words, they modify stability measures, e.g., by honing existing skills or developing feedback mechanisms, to improve configuration survivability against fluctuations. Active systems can also improve the survivability function, by adding skills, which is equivalent to adding terms in the survivability function, or by changing the environment, which is equivalent to changing the noise parameters. Biology is much richer because the noise , e.g., predators and resources, is also adaptive and changing in response. Biological reproduction also has an equivalent in passive systems - a line of descendents can be regarded as a sequence of the stable configurations in self-organised steady states. In short, as the objective of adaptation, evolution, reproduction, and feedback mechanisms is always to increase survivability of individuals and species against fluctuations and constraints, they can be described, in principle, by the proposed survivability statistical mechanics.
The ’in principle’ is important - to model complex organisms with the proposed formalism would require taking into consideration a large number of noise sources, functionalities, skills, etc. in the survivability function. Therefore, this formalism is practically applicable only to simple biological systems. It should be emphasised, however, that this is a quantitative difficulty rather than a conceptual one.
VIII. Concluding discussion
A general principle, which governs SO in out-of-equilibrium driven disordered systems, has been formulated. It states that SO emerges when a minute subset of system configurations are exceptionally stable and survive the noise, generated by the driving and other constraints, sufficiently long to dominate observations. The apparent reduction in entropy is because the overwhelming majority of theoretically possible configurations are dismantled by the noise and are too short-lived on observation timescales. This principle extends an earlier principle, based on energy considerations Cyetal21 ; Keetal23 , to SO phenomena determined by general mechanisms Jietal25 ; HeMo95 ; Chetal00 .
Based on this principle, a statistical mechanical-like formalism has been proposed to model the characteristics of the long-term self-organised state. It involves identification of quasi-particles, their DOFs, stability measures, and a survivability function, whose value decreases with the stability measures. A partition function has been constructed, in which the occurrence probabilities of system configurations are determined by Boltzmann-like exponentials to accommodate extensivity. The SO is determined by the high occurrence probability of configurations that minimise locally the survivability function.
The formalism makes possible derivation of relevant expectation values of the self-organised state, equivalents of equations of state, and phase-like transitions. This predictive power has been demonstrated for two examples: a cooperative SO in 2D granular experiments and the laning phenomenon. In the former, expectation values of local structural and stress stability measures have been derived, as well as the equivalent of an equation of state. The extension to three-dimensional granular systems has also been outlined and discussed in detail. In the second example, a simple model has been constructed for the laning phenomenon, which enabled a derivation of the mean lane speed. Surprisingly, the method provides a prediction of a non-trivial phase-like transition as a function of noise and bias parameters. In both these examples, the SO is not determined by energy considerations. While hardly any literature exists about relations between stability measures and long-term SO characteristics, as defined here, these are straightforward to obtain in simulations and experiments, which would help test some of the explicit results derived here.
@@ VIII. Concluding discussion
There are parallels between this formulation and its analogue in thermal systems, in which the microstates are the system-wide configurations, particles’ energies and potentials are used as stability measures, and the thermal survivability function is the free energy. Like the free energy, the survivability function depends on the quasi-particles’ DOFs and its lowering increases a microstate’s occurrence probability and hence its survivability. Following this rationale, some equilibrium macrostates can also be viewed as self-organised. An example is solidification under temperature lowering, where the overwhelming number of microstates with very energetic particles are strongly unstable and only the minute ordered microstates survive to be observed. The lowering of temperature is analogous to increasing $J_{i}$ in (4).
There are, however, differences. The energy in thermal systems is a constant of the motion, from which dynamic equations of motion of the DOFs can be derived. The survivability function only describes stability in long-term out-of-equilibrium self-organised steady states and, as far as this author knows, cannot play such a role. Additionally, thermal systems incur indigenous thermal fluctuations while in athermal systems it is the driving that gives rise to noise. Nevertheless, the similarities to traditional statistical mechanics are useful as guidelines for constructing general models of SO.
The principle and the modelling approach are applicable, in principle, to SO of biological systems. In this context, the principle translates to ‘the fittest survive’, i.e., the more stable the biological system the more fit it is to survive its noisy environment. The equivalence of stability measures and survivability functions between SO in such systems and in inanimate processes has been pointed out. As discussed, adaptability, reproduction, and feedback mechanisms can also be accommodated within the proposed formalism, as mechanisms by which biological systems modify the survivability function to improve stability measures and fitness. While the large number of DOFs, stability measures, and noise sources make it difficult to model complex organisms and bio-networks, there is a clear potential application to relatively simple ones.
The abundance of self-organised phenomena offers many opportunities to test this approach. The procedure is to identify quasi-particles that are sensitive the SO, identify stability measures that depend of the quasi-particles’ DOFs, and construct a survivability function that decreases with configurational stability. The dependence of the characteristics of the self-organised state on measurable parameters can also be included indirectly in the survivability function. For example, in the granular example, the inertial number can be controlled by shear rates, which affects the noise parameters. Similarly, the density should be correlated with the cell order distribution, $Q_{k}$. Including experimentally controlled parameters makes this formalism very flexible. The expectation values of characteristics of the long-term self-organised states, derived with the formalism can be then tested against observations. This author is looking forward to such tests.
@@ Acknowledgement:
Discussions with J.L. England and D.I. Goldman are gratefully acknowledged, as well as reviewers’ comments.